OpenStudy (ria23):

Can someone work this problem out for me so I can understand how to do it and finish this test? .3. What is the determinant of the matrix? [5 6] [2 9] (It's all one big bracket.)

OpenStudy (shiraz14):

Det = ad - bc = 45-12 = 33

OpenStudy (ria23):

:o oh. Thank yhu. ^.^

OpenStudy (ria23):

. . . I can't get the next one. The answer I got isn't listed. ; n;

OpenStudy (ria23):

What is the determinant of the matrix? [-4 -5] [2 -6] I got 24-(-30)= 54 My choices are 34, 14, -14, -34.

OpenStudy (ria23):

*Facepalm* Nevermind.

OpenStudy (shiraz14):

How did you get the -30?

OpenStudy (ria23):

xD I have no idea. I must have looked at that wrong.

OpenStudy (shiraz14):

It should be (-4*-6)-(2*-5) = 24-(-10) = 24+10 = 34

OpenStudy (ria23):

That's what I got. Thank yhu ^.^ I need help with 3 more questions, they're slightly different. Do yhu think yhu could walk me through how to do them?

OpenStudy (shiraz14):

Post them in a different thread. Thanks.

OpenStudy (ria23):

No problem. C: